# On an Erdős-Szekeres Game

### v2

@@ Abstract
We consider a 2-player permutation game inspired by the celebrated Erdős-Szekeres Theorem. The game depends on two positive integer parameters $a$ and $b$ and we determine the winner and give a winning strategy when $a\geq b$ and $b\in\left\{2,3,4,5\right\}$.
@@ 1 Introduction
Let $\mathcal{S}_{n}$ be the set of all permutations of $\{1,2,\dots,n\}$. We say that permutation $\pi\in\mathcal{S}_{n}$ contains $\rho\in\mathcal{S}_{m}$ if $\pi$ contains a subsequence $\pi_{i_{1}}\pi_{i_{2}}\cdots\pi_{i_{m}}$ such that $i_{1}<i_{2}<\cdots<i_{m}$ and $\pi_{i_{a}}<\pi_{i_{b}}$ if and only if $\rho_{a}<\rho_{b}$. Otherwise we say $\pi$ avoids $\rho$. One of the oldest theorems that can rephrased in terms of permutation patterns is the Erdős-Szekeres Theorem [2], which was first published in 1935, and is phrased in terms of patterns as Theorem 1.
@@
While there are a variety of proofs of 1, one of the most concise was given by Seidenberg [8] in 1959, using an application of the pigeonhole principle, which we revisit here:
@@
Let $\pi=\pi_{1}\cdots\pi_{n}\in\mathcal{S}_{n}$. For each $\pi_{i}$ associate an ordered pair $(a_{i},d_{i})$ where $a_{i}$ is the length of the longest increasing subsequence of $\pi$ ending in $\pi_{i}$ and $d_{i}$ is the length of the longest decreasing subsequence of $\pi$ ending in $\pi_{i}$. Clearly, for all $i$, $a_{i}\geq 1$ and $d_{i}\geq 1$ since the digit $\pi_{i}$ is itself an increasing (resp. decreasing) subsequence of length 1.
@@
However, we also have that if $i\neq j$, then $(a_{i},d_{i})\neq(a_{j},d_{j})$. This follows from the fact that the digits of $\pi$ are distinct members of $\left\{1,\dots,n\right\}$. Without loss of generality, suppose $i<j$. If $\pi_{i}<\pi_{j}$, then $a_{i}<a_{j}$ since appending $\pi_{j}$ onto the increasing subsequence of length $a_{i}$ ending at $\pi_{i}$ produces an increasing subsequence of length $a_{i}+1$ ending at $\pi_{j}$. Similarly, if $\pi_{i}>\pi_{j}$, then $d_{i}<d_{j}$.
@@
We have $n$ distinct ordered pairs of positive integers associated with $\pi$. If $a_{i}\geq a$ or $d_{i}\geq b$ for some $i$ then $\pi$ contains an $I_{a}$ or a $J_{b}$ pattern. So, if $\pi$ avoids $I_{a}$ and $J_{b}$, then $11\leq a_{i}\leq a-$ 1 and $11\leq d_{i}\leq b-$ 1 for all $i$, which means $n\leq(a-1)(b-1)$. Taking the contrapositive, if $n\geq(a-1)(b-1)+1$, then $\pi$ contains either $I_{a}$ or $J_{b}$ as a pattern.
@@
In 1983, Harary, Sagan, and West [3] studied a game based on Theorem 1 with rules as follows: Consider the set of integers $\left\{1,2,\dots,ab+1\right\}$. Two players take turns selecting numbers from this set until either an increasing subsequence of length $a+1$ or a decreasing subsequence of length $b+1$ is formed. In the achievement version of the game, the first player to complete such a subsequence wins. In the avoidance version of the game, the first player to complete such a subsequence loses. Their analysis is computer-aided, and is limited by the computer memory available at the time. In particular, each state of the game can be labeled as winning or losing for player 1 based on an analysis of subsequent possible moves. They determined the winning player for games where $ab+1\leq 15$, and since the tree of game states grows exponentially in $a$ and $b$, they predicted that it is prohibitive to push computer analysis much further. While they determined the winner of the game for particular small values of $a$ and $b$, they were unable to find a general winning strategy.
@@
In 2009, Albert et. al. [1] considered the game of Harary, Sagan, and West as well as a number of generalizations. In particular, they also considered the play of the game on $\mathbb{Q}$ rather than on a finite set. In this latter setting they determined that the winner of the achievement game on $\mathbb{Q}$ with parameters $a$ and $b$ is the same as the winner of the avoidance game with parameters $a-1$ and $b-1$ (Proposition 7 of [1]). They also determined the winner of the achievement game for $a\geq b$ and $b\leq 5$ (Theorem 9 of [1]). The cases where $b=2$ and $b=3$ are simple, but they gave an explicit strategy for a first player win in the cases where $b=4$ and $b=5$. In addition, for arbitrarily large $b$, they determine that the game where $a=b$ is a first player win for $b\geq 4$.
In 2021, the current author proposed a variation on this game as part of a public lecture [5]. Instead of proposing a 2-player competitive game, the new version is collaborative, and instead of picking specific digits from a set, players append a new digit onto a permutation prefix. In particular, after the first $n$ moves, the current game state is a permutation $\pi\in\mathcal{S}_{n}$, and the next player can play any digit in $\left\{1,\dots,n+1\right\}$ as their move. After the first move, the game is always in state $\pi=1$. However, on subsequent moves, if a player chooses $m\in\left\{1,\dots,n+1\right\}$, we let $A=\left\{i\middle|\pi_{i}\leq m-1\right\}$ and let $B=\left\{i\middle|\pi_{i}\geq m\right\}$ and the new game state becomes $\pi^{\prime}=\pi_{1}^{\prime}\pi_{2}^{\prime}\cdots\pi_{n}^{\prime}m\in\mathcal{S}_{n+1}$ where $\pi_{i}^{\prime}=\pi_{i}$ if $i\in A$ and $\pi_{i}^{\prime}=\pi_{i}+1$ if $i\in B$. Notice that $\pi_{1}^{\prime}\pi_{2}^{\prime}\cdots\pi_{n}^{\prime}$ forms a $\pi$ pattern in $\pi^{\prime}$, that is, the relative order of the initial digits is unchanged. This relative-order requirement is equivalent to playing a collaborative form of the game of [3] on $\mathbb{Q}$ and is thus a variation on the competitive game of [1]. The game ends when the most recent move completes an $I_{a}$ pattern or a $J_{b}$ pattern. The collaborative version of this game was generated as a teaching tool to build intuition about Theorem 1 for audience members who were unfamiliar with the theorem. Optimal collaborative game play achieves a permutation of length $(a-1)(b-1)+1$.
In [6], the current author analyzed a question that arose from continuing to use this collaborative game as a teaching tool: how many ways are there to play the game optimally? For example, if $a\geq b=2$, optimal game play is to build an increasing permutation of length $(a-1)(2-1)+1=a$, and there is only one such permutation. However, if $a\geq b>2$, there are more maximum length permutations. These permutations can be counted by mapping permutations avoiding $I_{a}$ and $J_{b}$ to pairs of same-shaped standard Young tableaux with at most $a-1$ columns and $b-1$ rows using the RSK algorithm, and then by using the hook length formula to count the number of such pairs. This process was elaborated on by Stanley [7], who cited work of Schensted [9] in a solution to a problem posed in American Mathematical Monthly in 1969. While this strategy answers the general enumeration question it is worth noting that the $a\geq b=3$ case has a particularly nice enumeration: there are $\left(C_{a-1}\right)^{2}$ permutations of length $2(a-1)$ avoiding both $I_{a}$ and $J_{3}$, where $C_{n}$ denotes the $n$th Catalan number, i.e., $C_{n}=\dfrac{\binom{2n}{n}}{n+1}$, and therefore there are $\left(C_{a-1}\right)^{2}\cdot a$ maximum length permutations in the $a\geq b=3$ game. The current author proved this known enumeration via simpler techniques, i.e., using a bijection between the $\left(C_{a-1}\right)^{2}$ permutations in question and pairs of parentheses arrangements. Of note, the bijection focused on tracking the positions and the values of the left-to-right maxima of the optimal play permutations.
@@
In this paper, we consider the 2-player competitive version of this game, which is a reformulation of the game on $\mathbb{Q}$ studied in [1]. In particular, as in the collaborative version, after the first $n$ moves, the current game state is a permutation of length $n$, and the next player can play any digit in $\left\{1,\dots,n+1\right\}$ as their move, appending it to the current permutation. We will primarily analyze the game where the first player who completes an $I_{a}$ pattern or a $J_{b}$ pattern loses. In contrast to the analysis of Harary, Sagan, and West’s game, where no general strategy was found, we give strategies that work for specific choices of $b$ but for any $a\geq b$. In contrast to the formulation in [1], we model the game as shading cells in a 2-dimensional grid while they model it by a building a ternary word with certain restrictions. For completeness, we show how to use the grid-shading model to see that the parity of $a$ determines the winner when $b=2$ and to give a strategy for a first player win in the avoidance game for $b=3$ and $b=4$. Since [1] focuses primarily on the attainment game, these strategies match their results for $b=3$, $b=4$, and $b=5$. In addition, we use our grid-shading model to give a winning strategy for the $b=5$ avoidance game, which goes beyond the scope of winning strategies given in [1].
@@
The organization of the rest of this paper is as follows. In Section 2 we develop a visual way of representing game moves, inspired by Seidenberg’s proof of Theorem 1, which was given above. In Sections 3, 4, 5, and 6 we give a general winning strategy for the cases where $a\geq b$ and $2\leq b\leq 5$. In Section 7 we consider the version of the game where the first player who completes an $I_{a}$ pattern or a $J_{b}$ pattern wins the game. Finally, we conclude with topics for future investigation.
@@ 2 Representation of Moves
As described in Section 1, we consider a 2-player game where players take turns appending a new digit onto a permutation. On the $n$th turn of the game, a player may play any number in $\left\{1,\dots,n\right\}$. After the $n$th turn of the game, the current game state is a permutation $\pi$ of length $n$. For all $i\geq 1$, the pattern formed by the first $i$ digits remains unchanged as the game progresses, but we need to track increasing and decreasing subsequences within the permutation being constructed. We call a game that ends when a player completes an $I_{a}$ or a $J_{b}$ pattern an $(a,b)$-game. The Seidenberg proof of Theorem 1 tracks increasing and decreasing subsequences using ordered pairs of positive integers. Motivated by this representation we define the board of an $(a,b)$-game to be a grid of cell cells with $b-1$ rows and $a-1$ columns. Each cell is indexed by the ordered pair $(c,r)$ where $c$ denotes the column number and $r$ denotes the row number of the cell.
Now, for each move of an $(a,b)$-game, we shade a cell $(c,r)$ if the longest increasing subsequence of $\pi$ ending in $\pi_{n}$ has length $c$ and the longest decreasing subsequence of $\pi$ ending in $\pi_{n}$ has length $r$. Figure 1 shows the board in the $(6,5)$-game corresponding to $\pi=163425$. To check, $\pi_{1}=1$ corresponds to $(1,1)$, $\pi_{2}=6$ corresponds to $(2,1)$, $\pi_{3}=3$ corresponds to $(2,2)$, $\pi_{4}=4$ corresponds to $(3,2)$, $\pi_{5}=2$ corresponds to $(2,3)$, and $\pi_{6}=5$ corresponds to $(4,2)$. Note that the increasing subsequence length is given first for consistency of notation with the name of $(a,b)$-game, while these values correspond to column numbers (rather than row numbers) in the board for vertically efficient use of the page throughout this manuscript.
At the $n$th move, we know that for each $j<n$ either $\pi_{j}<\pi_{n}$ (in which case the column number of the $n$th shaded cell is larger than the column number of the $j$th shaded cell) or $\pi_{j}>\pi_{n}$ (in which case the row number of the $n$th shaded cell is larger than the row number of the $j$th shaded cell). So, each newly shaded cell must be pairwise further to the right or further below each previously shaded cell. To this end, in addition to the shaded cells we call a cell $(c^{*},r^{*})$ eliminated if $(c,r)$ is shaded and both $c^{*}\leq c$ and $r^{*}\leq r$. The cells $(3,1)$, $(4,1)$, $(1,2)$, and $(1,3)$ are eliminated in Figure 1 and thus marked with hatching. These are cells that are ineligible to become shaded in future turns, based on the permutation formed so far in the game. The shaded cells are a subset of the eliminated cells at any point of the game.
@@
We now make some observations about the shaded region of a board at any point in an $(a,b)$-game.
First, by definition of eliminated cells, at any point in the game, the set of shaded and eliminated board cells forms one contiguous region. Moreover, this region tells us clearly which cells are available to be shaded by the next move of the game. We refer to cells that are not shaded and not eliminated as open cells.
@@
Suppose $S$ is the set of shaded and eliminated cells after the $n$th turn of the $(a,b)$-game. Then move $n+1$ corresponds to shading an open cell that is edge-adjacent to a cell is $S$. Moreover, every open cell that is edge-adjacent to $S$ corresponds to a possible next move.
@@
First, consider the extreme cases where $\pi_{n+1}=1$ or $\pi_{n+1}=n+1$. If $\pi_{n+1}=1$, then playing $\pi_{n+1}$ corresponds to cell $(1,d+1)$. This is edge adjacent to $S$ since some cell in $S$ has row number $d$, and therefore cell $(1,d)$ is either shaded or eliminated. Similarly, if $\pi_{n+1}=n+1$, playing $\pi_{n+1}$ corresponds to cell $(i+1,1)$. This is also edge adjacent to $S$ since some cell has column number $i$, and therefore cell $(i,1)$ is either shaded or eliminated.
Now suppose that $\pi_{n+1}=j$ ($11<j<n+$ 1) corresponds to shading the cell $(c,r)$. We consider what cell would be shaded if $\pi_{n+1}={j+1}$ instead in four cases. In all cases, the digits $\pi_{1}\cdots\pi_{n}$ have the same relative order. The only difference is that if $\pi_{n+1}=j$, there exists $1\leq k\leq n$ such that $\pi_{k}=j+1$, and now when $\pi_{n+1}=j+1$, instead $\pi_{k}=j$.
  • $\pi_{k}$ played no role in the increasing subsequence of length $c$ or the decreasing subsequence of length $r$ ending in $\pi_{n+1}=j$, and so $\pi_{n+1}=j$ and $\pi_{n+1}=j+1$ result in shading the same $(c,r)$ cell.
  • $\pi_{k}$ corresponds to shading cell $(c,r-1)$ and $\pi_{k}$ was part of the decreasing subsequence of length $r$ when $\pi_{n+1}=j$. In the situation where $\pi_{n+1}=j+1$, $\pi_{k}\pi_{n+1}$ are the last two digits in an increasing subsequence of length $c+1$, while $\pi_{n+1}$ ends a decreasing subsequence of length $r-1$. In other words, $\pi_{n+1}=j+1$ corresponds to shading $(c+1,r-1)$.
  • $\pi_{k}$ corresponds to shading cell $(c^{*},r-1)$ where $c^{*}<c$ and $\pi_{k}=j+1$ played a role in forming the decreasing subsequence of length $r$ when $\pi_{n+1}=j$. Now that $\pi_{n+1}={j+1}$, $\pi_{n+1}$ completes a decreasing subsequence of length $r-1$, while $c$ remains unchanged. In other words, $\pi_{n+1}=j+1$ corresponds to shading $(c,r-1)$.
@@
These four cases describe the set of possible cells that can be shaded by various choices of $\pi_{n+1}$. We gave specific examples that showed $(1,d+1)$ and $(i+1,1)$ are possible. We also saw that each open cell that could be shaded by a choice of $\pi_{n+1}$ differs in row and/or column number by at most 1 from another open cell. In fact, the only time when both numbers change is when they are immediately below and immediately right of a shaded cell that forms a southeast corner of $S$. This uniquely describes the open cells that are edge adjacent to $S$.
As an example, consider Figure 2, which shows the possible ordered pairs corresponding to next moves in a $(6,5)$-game whose current permutation is $\pi=163425$. The left side of the figure shows the plots of the points $(i,\pi_{i})$ with ordered pairs given for various choices of $\pi_{7}$, while the right side of the figure shows the shaded and eliminated cells after the sixth turn, and highlights the cell corresponding to choices of $\pi_{7}$.
$(5,1)$$(5,2)$$(4,3)$$(3,3)$$(3,3)$$(2,4)$$(1,4)$
@@
Finally, the player who shades the $(a-1,b-1)$ cell is the winner since this cell being shaded means that the entire board is either shaded or eliminated, and the next player must add a digit to the permutation that either completes an $I_{a}$ pattern or a $J_{b}$ pattern. Now that we have made this translation from digits to boards, we may play the Erdős-Szekeres game as a game of shading cells on a board, rather than thinking merely in terms of one-line permutation notation.
@@
  • For each subsequent move, a player shades a cell that is edge-adjacent to the shaded region. All cells that are above and/or right of their chosen cell should also be shaded (i.e. eliminated).
@@
Notice that there are many permutations that may correspond to the same shaded board. As a small example, both $\pi=132$ and $\pi=312$ correspond to a board where a $2\times 2$ region has been shaded. However, these are both permutations where the most recent digit corresponds to the label $(2,2)$, and so the same amount of progress has been made towards forming an $I_{a}$ or $J_{b}$ pattern. In terms of tracking a win or loss in the permutation game, no information has been lost.
Finally, we remark that this region of shaded and eliminated cells encodes exactly the same information about $\pi$ as the ternary words used in [1]. The notation in Albert et. al. is motivated directly by Schensted’s bumping algorithm in [9] rather than Seidenberg’s proof. They represent game states as words on the symbols R, B, and P with at least one P, not starting with B, not ending with R, and not containing RB as a factor. Notice that the boundary of the shaded/eliminated region in our representation is contiguous. If we consider shaded and eliminated cells with bottom and/or right edges on the boundary of the shaded region, we can traverse these cells from lower left to upper right, and record R if only the bottom edge is on the boundary of the shaded region, B if only the right edge is on the boundary of the shaded region, and P if both the bottom and right edges are on the boundary of the shaded region. For example, the gray shaded and eliminated region corresponding to $\pi=163425$ shown in Figure 1 corresponds to the word $RPRPB$. The word notation of Albert et. al. has the advantage that it is language theoretic. The grid presentation in this paper has the advantage that one can play the game focused on geometric information, which can make it easier to quickly ascertain the state of the game without counting or computation.
In the following sections, we articulate a strategy for winning this permutation game in terms of board shading. The strategies for $2\leq b\leq 4$ exactly match the strategies that were presented in language theoretic terms in [1] while the strategy for $b=5$ is completely new and extends their results.

### v3

@@ Abstract
We consider a two-player permutation game inspired by the celebrated Erdős-Szekeres Theorem. The game depends on two positive integer parameters $a$ and $b$ and we determine the winner and give a winning strategy when $a\geq b$ and $b\in\left\{2,3,4,5\right\}$.
@@ 1 Introduction
Let $\mathcal{S}_{n}$ be the set of all permutations of $\{1,2,\dots,n\}$. We say that permutation $\pi\in\mathcal{S}_{n}$ contains $\rho\in\mathcal{S}_{m}$ if $\pi$ contains a subsequence $\pi_{i_{1}}\pi_{i_{2}}\cdots\pi_{i_{m}}$ such that $i_{1}<i_{2}<\cdots<i_{m}$ and $\pi_{i_{a}}<\pi_{i_{b}}$ if and only if $\rho_{a}<\rho_{b}$. Otherwise we say $\pi$ avoids $\rho$. One of the oldest theorems that can rephrased in terms of permutation patterns is the Erdős-Szekeres Theorem [2], which was first published in 1935, and is phrased in terms of patterns as Theorem 1 below.
@@
While there are a variety of proofs of Theorem 1, one of the most concise was given by Seidenberg [7] in 1959, using an application of the pigeonhole principle, which we revisit here:
@@
Let $\pi=\pi_{1}\cdots\pi_{n}\in\mathcal{S}_{n}$. For each $\pi_{i}$ associate an ordered pair $(a_{i},d_{i})$ where $a_{i}$ is the length of the longest increasing subsequence of $\pi$ ending in $\pi_{i}$ and $d_{i}$ is the length of the longest decreasing subsequence of $\pi$ ending in $\pi_{i}$. Clearly, for all $i$, $a_{i}\geq 1$ and $d_{i}\geq 1$ since the entry $\pi_{i}$ is itself an increasing (resp. decreasing) subsequence of length 1.
@@
However, we also have that if $i\neq j$, then $(a_{i},d_{i})\neq(a_{j},d_{j})$. This follows from the fact that the entries of $\pi$ are distinct members of $\left\{1,\dots,n\right\}$. Without loss of generality, suppose $i<j$. If $\pi_{i}<\pi_{j}$, then $a_{i}<a_{j}$ since appending $\pi_{j}$ onto the increasing subsequence of length $a_{i}$ ending at $\pi_{i}$ produces an increasing subsequence of length $a_{i}+1$ ending at $\pi_{j}$. Similarly, if $\pi_{i}>\pi_{j}$, then $d_{i}<d_{j}$.
@@
We have $n$ distinct ordered pairs of positive integers associated with $\pi$. If $a_{i}\geq a$ or $d_{i}\geq b$ for some $i$ then $\pi$ contains an $I_{a}$ or a $J_{b}$ pattern. So, if $\pi$ avoids $I_{a}$ and $J_{b}$, then $1\leq a_{i}\leq a-1$ and $1\leq d_{i}\leq b-1$ for all $i$, which means $n\leq(a-1)(b-1)$. Taking the contrapositive, if $n\geq(a-1)(b-1)+1$, then $\pi$ contains either $I_{a}$ or $J_{b}$ as a pattern.
@@
In 1983, Harary, Sagan, and West [3] studied a game based on Theorem 1 with rules as follows: Consider the set of integers $\left\{1,2,\dots,(a-1)(b-1)+1\right\}$. Two players take turns selecting numbers from this set until either an increasing subsequence of length $a$ or a decreasing subsequence of length $b$ is formed. In normal play, the first player to complete such a subsequence wins. In the misère version of the game, the first player to complete such a subsequence loses. The analysis in [3] is computer-aided, and is limited by the computer memory available at the time. In particular, each state of the game can be labeled as winning or losing for player 1 based on an analysis of subsequent possible moves. They determined the winning player for games where $(a-1)(b-1)+1\leq 15$, and since the tree of game states grows exponentially in $a$ and $b$, they predicted that it is prohibitive to push computer analysis much further.
@@
In 2009, Albert et. al. [1] considered the game of Harary, Sagan, and West as well as a number of generalizations. In particular, they also considered the play of the game on $\mathbb{Q}$ rather than on a finite set. In this latter setting they determined that the winner of the game on $\mathbb{Q}$ with parameters $a$ and $b$ is the same as the winner of the misère game with parameters a1a-1 and b1b-1 (Proposition 7 of [1]). They also determined the winner of the game for $a\geq b$ and $b\leq 5$ (Theorem 9 of [1]). The cases where $b=2$ and $b=3$ are simple, but they gave an explicit strategy for a first player win in the cases where $b=4$ and $b=5$. In addition, for arbitrarily large $b$, they determine that the game where $a=b$ is a first player win for $b\geq 4$.
A collaborative version of this game has also been used as a teaching tool to build intuition about Theorem 1 (see [5]). Of note, the current author [6] determined how many maximum length permutations avoid $I_{a}$ and $J_{b}$ for the $a\geq b=3$ case via a bijection that tracks the positions and the values of the left-to-right maxima of the relevant permutations.
@@
In this paper, we consider the two-player competitive version of this game, which is a reformulation of the game on $\mathbb{Q}$ studied in [1]. In particular, after the first $n$ moves, the current game state is a permutation of length $n$, and the next player can play any entry in $\left\{1,\dots,n+1\right\}$ as their move, appending it to the current permutation. We will primarily analyze the misère game, where the first player who completes an $I_{a}$ pattern or a $J_{b}$ pattern loses. We give strategies that work for specific choices of $b$ but for any $a\geq b$. In contrast to the formulation in [1], we model the game as shading cells in a 2-dimensional grid while they model it by a building a ternary word with certain restrictions. For completeness, we show how to use the grid-shading model to see that the parity of $a$ determines the winner when $b=2$ and to give a strategy for a first player win in the avoidance game for $b=3$ and $b=4$. Since [1] focuses primarily on normal play, these strategies match their results for $b=3$, $b=4$, and $b=5$. In addition, we use our grid-shading model to give a winning strategy for the $b=5$ case of misère play, which goes beyond the scope of winning strategies given in [1].
@@
The organization of the rest of this paper is as follows. In Section 2 we develop a visual way of representing game moves, inspired by Seidenberg’s proof of Theorem 1, which was given above. In Sections 3, 4, 5, and 6 we give a general winning strategy for the cases of misère play where $a\geq b$ and $2\leq b\leq 5$. Finally, in Section 7, we conclude with observations and topics for future investigation.
@@ 2 Representation of Moves
As described in Section 1, we consider a two-player game where players take turns appending a new entry onto a permutation. On the $n$th turn of the game, a player may play any number in $\left\{1,\dots,n\right\}$. After the $n$th turn of the game, the current game state is a permutation $\pi$ of length $n$. For all $i\geq 1$, the pattern formed by the first $i$ entries remains unchanged as the game progresses, but we need to track increasing and decreasing subsequences within the permutation being constructed. We call a game that ends when a player completes an $I_{a}$ or a $J_{b}$ pattern an $(a,b)$-game. The Seidenberg proof of Theorem 1 tracks increasing and decreasing subsequences using ordered pairs of positive integers. Motivated by this representation we define the board of an $(a,b)$-game to be a grid with a1a-1 columns and b1b-1 rows of cells. Each cell is indexed by the ordered pair $(c,r)$ where $c$ denotes the column number and $r$ denotes the row number of the cell.
Now, for the $n$th move of an $(a,b)$-game, we shade a cell $(c,r)$ if the longest increasing subsequence of $\pi$ ending in $\pi_{n}$ has length $c$ and the longest decreasing subsequence of $\pi$ ending in $\pi_{n}$ has length $r$. Figure 1 shows the board in the $(6,5)$-game corresponding to $\pi=163425$. (Note that the hatched regions in the figure are not shaded and will be introduced later.) To check, $\pi_{1}=1$ corresponds to $(1,1)$, $\pi_{2}=6$ corresponds to $(2,1)$, $\pi_{3}=3$ corresponds to $(2,2)$, $\pi_{4}=4$ corresponds to $(3,2)$, $\pi_{5}=2$ corresponds to $(2,3)$, and $\pi_{6}=5$ corresponds to $(4,2)$. Note that the increasing subsequence length is given first for consistency of notation with the name of $(a,b)$-game, while these values correspond to column numbers (rather than row numbers) in the board for vertically efficient use of the page throughout this manuscript.
At the $n$th move, we know that for each $j<n$ either $\pi_{j}<\pi_{n}$ (in which case the column number of the $n$th shaded cell is larger than the column number of the $j$th shaded cell) or $\pi_{j}>\pi_{n}$ (in which case the row number of the $n$th shaded cell is larger than the row number of the $j$th shaded cell). To this end, in addition to the shaded cells we call a cell $(c^{*},r^{*})$ eliminated if $(c,r)$ is shaded and both $c^{*}\leq c$ and $r^{*}\leq r$. The cells $(3,1)$, $(4,1)$, $(1,2)$, and $(1,3)$ are eliminated in Figure 1 and thus marked with hatching. These are cells that are ineligible to become shaded in future turns, based on the permutation formed so far in the game. The shaded cells are a subset of the eliminated cells at any point of the game.
@@
We now make some observations about the eliminated region of a board at any point in an $(a,b)$-game.
First, by definition of eliminated cells, at any point in the game, the set of eliminated board cells forms one contiguous region. Moreover, this region tells us clearly which cells are available to be shaded by the next move of the game. We refer to cells that are not eliminated as open cells.
@@
Suppose $S$ is the set of eliminated cells after the $n$th turn of the $(a,b)$-game. Then move $n+1$ corresponds to shading an open cell that is edge-adjacent to a cell is $S$. Moreover, every open cell that is edge-adjacent to $S$ corresponds to a possible next move.
@@
First, consider the extreme cases where $\pi_{n+1}=1$ or $\pi_{n+1}=n+1$. If $\pi_{n+1}=1$, then playing $\pi_{n+1}$ corresponds to cell $(1,d+1)$. This is edge adjacent to $S$ since some cell in $S$ has row number $d$, and therefore cell $(1,d)$ is eliminated. Similarly, if $\pi_{n+1}=n+1$, playing $\pi_{n+1}$ corresponds to cell $(i+1,1)$. This is also edge adjacent to $S$ since some cell has column number $i$, and therefore cell $(i,1)$ is eliminated.
Now suppose that $\pi_{n+1}=j$ ($1<j<n+1$) corresponds to shading the cell $(c,r)$. We consider what cell would be shaded if $\pi_{n+1}={j+1}$ instead in four cases. In all cases, the entries $\pi_{1}\cdots\pi_{n}$ have the same relative order. The only difference is that if $\pi_{n+1}=j$, there exists $1\leq k\leq n$ such that $\pi_{k}=j+1$, while if $\pi_{n+1}=j+1$, instead $\pi_{k}=j$. When $\pi_{k}=j+1$ and $\pi_{n+1}=j$, it is not possible for $\pi_{k}$ to be part of the increasing subsequence of length $c$ ending in $\pi_{n+1}$ since $\pi_{k}>\pi_{n+1}$.
  • $\pi_{k}$ played no role in the decreasing subsequence of length $r$ ending in $\pi_{n+1}=j$, and so $\pi_{n+1}=j$ and $\pi_{n+1}=j+1$ result in shading the same $(c,r)$ cell.
  • $\pi_{k}$ corresponds to shading cell (c,r1)(c,r-1) and $\pi_{k}$ was part of the decreasing subsequence of length $r$ when $\pi_{n+1}=j$. In the situation where $\pi_{n+1}=j+1$, $\pi_{k}\pi_{n+1}$ are the last two entries in an increasing subsequence of length $c+1$, while $\pi_{n+1}$ ends a decreasing subsequence of length r1r-1. In other words, $\pi_{n+1}=j+1$ corresponds to shading (c+1,r1)(c+1,r-1).
  • $\pi_{k}$ corresponds to shading cell $(c^{*},r-1)$ where $c^{*}<c$ and $\pi_{k}=j+1$ played a role in forming the decreasing subsequence of length $r$ when $\pi_{n+1}=j$. Now that $\pi_{n+1}={j+1}$, $\pi_{n+1}$ completes a decreasing subsequence of length r1r-1, while $c$ remains unchanged. In other words, $\pi_{n+1}=j+1$ corresponds to shading (c,r1)(c,r-1).
@@
These four cases describe the set of possible cells that can be shaded by various choices of $\pi_{n+1}$. We gave specific examples that showed $(1,d+1)$ and $(i+1,1)$ are possible. We also saw that each open cell that could be shaded by a choice of $\pi_{n+1}$ differs in row and/or column number by at most 1 from another open cell. In fact, the only time when both numbers change is when they are immediately below and immediately right of a eliminated cell that forms a southeast corner of $S$. This uniquely describes the open cells that are edge adjacent to $S$.
As an example, consider Figure 2, which shows the possible ordered pairs corresponding to next moves in a $(6,5)$-game whose current permutation is $\pi=163425$. The left side of the figure shows the the points $(i,\pi_{i})$ for $1\leq i\leq 6$ with ordered pairs given for various choices of $\pi_{7}$, while the right side of the figure shows the eliminated cells after the sixth turn, and highlights the cells corresponding to choices of $\pi_{7}$.
(5,1)(5,1)(5,2)(5,2)(4,3)(4,3)(3,3)(3,3)(3,3)(3,3)(2,4)(2,4)$(1,4)$
@@
Finally, the player who shades the (a1,b1)(a-1,b-1) cell is the winner since this cell being shaded means that the entire board is eliminated, and the next player must add a entry to the permutation that either completes an $I_{a}$ pattern or a $J_{b}$ pattern. Now that we have made this translation from entries to boards, we may play the Erdős-Szekeres game as a game of shading cells on a board, rather than thinking merely in terms of one-line permutation notation.
@@
  • For each subsequent move, a player shades a cell that is edge-adjacent to the eliminated region. All cells that are above and/or left of their chosen cell should also be eliminated.
@@
We remark that this game is effectively a variant of the classic combinatorial game Chomp.
Notice that there are many permutations that may correspond to eliminating the same region of a game board. As a small example, both $\pi=132$ and $\pi=312$ correspond to a board where a $2\times 2$ region has been eliminated. However, these are both permutations where the most recent entry corresponds to the label $(2,2)$, and so the same amount of progress has been made towards forming an $I_{a}$ or $J_{b}$ pattern. In terms of tracking a win or loss in the permutation game, no information has been lost.
Finally, we remark that this region of eliminated cells encodes exactly the same information about $\pi$ as the ternary words used in [1]. The notation in Albert et. al. is motivated directly by Schensted’s bumping algorithm in [8] rather than Seidenberg’s proof. They represent game states as words on the symbols R, B, and P with at least one P, not starting with B, not ending with R, and not containing RB as a factor. Notice that the boundary of the eliminated region in our representation is contiguous. If we consider eliminated cells with bottom and/or right edges on the boundary of the eliminated region, we can traverse these cells from lower left to upper right, and record R if only the bottom edge is on the boundary of the eliminated region, B if only the right edge is on the boundary of the eliminated region, and P if both the bottom and right edges are on the boundary of the shaded region. For example, the gray eliminated region corresponding to $\pi=163425$ shown in Figure 1 corresponds to the word $RPRPB$. The word notation of Albert et. al. has the advantage that it is language theoretic. The grid presentation in this paper has the advantage that one can play the game focused on visual geometric information.
In the following sections, we articulate a strategy for winning this permutation game in terms of board shading. The strategies for $2\leq b\leq 4$ exactly match the strategies that were presented in language theoretic terms in [1] while the strategy for $b=5$ is completely new and extends their results. We will show that the empty board is in class $\mathcal{N}$ (i.e., a next-player winning position) when $3\leq b\leq 5$. When $b=2$, whether the empty board is in class $\mathcal{N}$ or class $\mathcal{P}$ (i.e. a next-player losing position) depends on the parity of $a$.